# Bridging the Visual Gap: Fine-Tuning Multimodal Models with Knowledge-Adapted Captions

### v1

@@ Abstract
Moran Yanuka1       Assaf Ben Kish1        Yonatan Bitton2       Idan Szpektor2       Raja Giryes1 1Tel Aviv University       2Google Research
Figure 1: KnowAda identifies knowledge gaps of a VLM and adapts the dense caption accordingly. The KnowAda dense captions are better suited for downstream training of the VLM.
@@ 1 Introduction
Fine-tuning pretrained multimodal models for generating dense image captions is common in both research and practical applications, such as assisting visually impaired individuals. Recent work has focused on creating high-quality, descriptive captions through human annotations Onoe et al. (2024); Garg et al. (2024); Deitke et al. (2024) and synthetic generations from models like GPT-4 Chen et al. (2023a) and Gemini Singla et al. (2024). These datasets help fine-tune models to produce detailed descriptions in specific styles, differing from the zero-shot captions of pretrained models. However, smaller multimodal models (e.g., up to 7 billion parameters), crucial for real-time applications, often struggle with capturing visual details during fine-tuning, leading to hallucinations.
@@
Recent works on large language models (LLMs) have shown that fine-tuning primarily adapts pre-existing factual knowledge for specific tasks, with the majority of this knowledge being encoded during the pretraining phase Geva et al. (2020); Meng et al. (2022); Roberts et al. (2020). Rather than acquiring new information, fine-tuning typically activates and refines pretrained knowledge, which remains largely stable throughout the process Zhou et al. (2024b). Furthermore, Gekhman et al. (2024) demonstrated that fine-tuning on content not grounded in a model’s pre-existing factual-knowledge can lead to an increase in hallucinations. Motivated by these findings, we demonstrate that similar challenges extend beyond factual knowledge, affecting multimodal models when fine-tuned on information that is not rooted in their pretrained visual understanding.
@@
To evaluate dense image captions, two factors are crucial: descriptiveness, which captures the image’s details, and hallucination rate, which measures factual accuracy. Traditional metrics, which focus on similarity to reference captions Papineni et al. (2002); Banerjee and Lavie (2005); Zhang et al. (2019); Reimers (2019) or CLIP-based alignment Sarto et al. (2023); Radford et al. (2021), struggle with long captions. These metrics penalize valid variations in phrasing and miss the distinction between factual accuracy and token overlap or semantic similarity, making them inadequate for detecting hallucinations in detailed captions. Existing hallucination metrics focus on short captions or object-level errors Rohrbach et al. (2018); Ben-Kish et al. (2023). In response, we propose Decomposed NLI (DNLI) a new evaluation framework that decomposes captions into propositions and measures their entailment with the ground truth description. DNLI provides a more reliable assessment of both descriptiveness and accuracy and highly aligns with human scores.
@@
Figure 3: Dense captioning descriptiveness precision-recall results for LLaVA-7B fine-tuned with DOCCI captions, adapted using different methods. “Trimmed” refers to naive removal of sentences, while “Gemini” involves prompting Gemini to simplify the caption by removing difficult details of varying degrees. KnowAda consistently achieves the best precision-recall balance. The original dataset corresponds to KnowAda with a threshold of $T=100\%$, where no information is classified as unknown.
@@ 2 KnowAda
We start by describing our caption adaptation method for dense captioning datasets. KnowAda is composed of three stages, as depicted in Figure 2. First, we generate visual questions from each dense caption using an LLM. Then, we probe the VLM’s pretrained visual knowledge and capabilities by finding the parts of the image caption the model has difficulty with. Then, an LLM adapts the image descriptions by editing-out unknown parts. We next elaborate on each step of the pipeline.
@@ 2.1 VLM Knowledge Probing
4. Calculate the difficulty of each question $q_{i}$ based on the ratio of correct to incorrect answers. The difficulty score $df_{i}$ for question $q_{i}$ is given by:
$df_{i}=\frac{|I_{i}|}{|C_{i}|+|I_{i}|}.$
@@
For example, consider the two questions shown in Figure 2 (ii). If the model correctly answered ‘‘Monorail’’ for the question ‘‘What kind of vehicle is visible in the image?’’ in 6 out of 10 sampled instances, with 4 incorrect answers, , the difficulty for this question is calculated as $df=\frac{4}{10}$ For the second question, ‘‘How many limousines are parked underneath the track?’’, the model correctly answered ‘‘3’’ only 4 out of 10 times, resulting in a difficulty of $df=\frac{6}{10}$. With a threshold of $T=50\%$, the first question is classified as known, while the second as unknown. However, if the threshold is raised to $T=70\%$, both questions would be considered unknown.
@@
2.2 Dataset Analysis
We begin by analyzing the characteristics of the dense captioning datasets and the effects of KnowAda on the resulting adapted captions.
@@ 3 Decomposed NLI Evaluation
Next, we introduce our proposed DNLI, a proposition-decomposition based evaluation framework for dense-captioning. It assesses the quality of a generated dense image caption using two key criteria: descriptiveness, which measures the accuracy and level of detail provided, and contradiction, which evaluates the extent to which the caption contradicts with the ground truth description. Drawing inspiration from prior work on paragraph summarization evaluation Ernst et al. (2021); Zhang and Bansal (2021); Ernst et al. (2021) and retrieval Chen et al. (2023b), DNLI assesses descriptiveness and consistency through proposition extraction, as outlined below and shown in Figure 4. For additional qualitative examples refer to Figure 10 in the appendix.
Propositional decomposition. Given a generated image description, we use a LLM to decompose it into a set of atomic propositions, capturing individual, verifiable claims. This enables fine-grained evaluation. To avoid duplicates, we instruct the model to include only unique propositions.
Natural Language Inference (NLI) analysis. We next evaluate the entailment of propositions generated from the image using both our automatic method, where a large language model (LLM) computes textual entailment, and through human annotators who directly assess visual entailment:
@@
Table 2: Statistics of the original and KnowAda adapted captions with $T=20\%$ over the original (human) and synthetic captions. $\bar{C}_{o}$ and $\bar{C}_{k}$ denotes the mean number of words in the original caption and caption after applying KnowAda, and $\bar{Q}_{unk}$ denotes the mean number of questions unknown to the model.
@@
Method Descriptiveness $\uparrow$ Contradiction $\downarrow$ # Words
@@
KnowAda Random 57.9 23.9 34.7 17.2 55
KnowAda 58.9 25.8 33.7 16.7 55
@@ 4 Experimental Settings
(i) Caption Trimming: A progressive method that removes varying numbers of sentences to simplify captions.
(ii) Gemini Simplification: An approach prompting Gemini to remove difficult details from captions to varying degrees.
@@ 5.3 Textual Entailment is Highly Correlated to Human Annotated Visual Entailment
Note that the results from our human annotation process revealed an average majority agreement of 82.6% for the “Contradicted” labels and 89.7% for the “Entailed” labels, indicating a high level of consensus among annotators.
5.4 Removing Unknown vs Random Information Ablation
@@ 7 Related Work
Dense Image Captioning  Traditional image captioning datasets like COCO Chen et al. (2015) often feature brief captions with limited visual detail. Recently, interest has grown in creating datasets with longer, more complex captions. For example, Onoe et al. 2024 and Garg et al., 2024 developed a dataset of 15K human-annotated, richly descriptive captions. Additionally, Chen et al., 2023a and Singla et al., 2024 employed GPT-4 and Gemini to generate detailed captions. We show that these datasets are more effective for fine-tuning when adapted to the specific model being trained.
@@
Training On Unknown Information. Several works examine the relationship between unknown information and downstream model performance. Gekhman et al. (2024) show that fine-tuning LLMs on low-confidence examples encourages hallucinations, and suggest that fine-tuning should not introduce new knowledge, but only teach the model to make use of existing knowlegde. Zhang et al. (2024b) suggest optimizing for higher-confidence responses via Direct Preference Optimization (DPO), where the confidence of each response is estimated via self-evaluation. Piché et al. (2024) train a utility function that encourages predicting only high-certainty responses. The function is trained over synthetic data, which is collected iteratively, and is composed of fused high-confidence predicted facts. Additionally, Xu et al. (2024) introduce a refusal mechanism which encourages the model to reject questions that do not align with its existing knowledge. We demonstrate that training a model on questions it is fundamentally lacking the ability to answer, rather than solely on factual knowledge it lacks—particularly in the multimodal domain—results in similar performance degradation. Moreover, our method only rejects the unknown parts of a data sample, while maintaining the information that is useful for training, thus avoiding throwing out useful training samples which only include some unknown parts.
@@
Data Curation in Fine-Tuning. Recent research has increasingly focused on refining the fine-tuning process for pretrained models, particularly through instruction tuning. Zhou et al. (2024b) showed that fine-tuning on a small but high-quality instruction tuning dataset can yield superior results. More closely related to our work, Li et al. (2023) explored training a language model on a subset of instruction tuning data, measuring loss discrepancies between samples with and without instructions to inform filtering strategies; however, this approach cannot be extended to general image captioning, where there are no instructions. Lin et al. (2024) proposed a method using a reference model to reduce the influence of noisy tokens by assigning them lower weights, while Chen et al. (2024) trained a filtering network to discard instruction tuning samples with consistently low loss. In contrast, our approach adapts specific parts of a caption that are misaligned with the model’s capabilities, rather than filtering or reweighing the entire caption. Importantly, our strategy leverages visual knowledge and focuses on sequences and conceptual coherence, moving beyond token-level adjustments.
@@ 8 Limitations
Secondly, although we demonstrate that KnowAda achieves the best performance in balancing the descriptiveness-hallucination trade-off, it does not completely resolve this issue. Reducing contradictions to the ground truth often results in a decrease in the descriptiveness recall (but not precision).
@@
9 Ethics Statement
This work focuses on measuring and mitigating hallucinations in visual-language models (VLMs). As such it is expected to increase the reliability of VLMs and the ability to measure their performance, which is important when using them in real world systems. This is expected to have a positive impact on the use of VLMs in the society. However, we do recognize that the foundation models used in the KnowAda construction and evaluation pipeline could propagate biases. We anticipate further research into such biases before relying on our work beyond the research environment. The human study performed in this work received the required IRB approval from our institution ethics committee.
@@ Appendix A Appendix
In this appendix, we describe the implementations details in Appendix A.1, qualitative examples in Appendix A.2, additional human annotations details in Appendix A.3 and additional experiments in Appendix A.4.
@@ A.1 Implementation Details
Proposition Entailment Evaluation Implementation Details  We use gemini-1.5-flash-001 both for the proposition extraction, as well as the textual entailment task. We constrain the output to be in JSON format, following the proposition extractin prompt shown in Figure 16 and textual entailment prompt in Figure 17.
@@ A.4 Additional Experiments
KnowAda Robustness Analysis. To assess the robustness of KnowAda, we follow the approach of Honovich et al. (2021); Wang et al. (2020), evaluating the percentage of unknown information successfully removed from the dense captions during KnowAda rewriting, as well as the percentage of known information successfully retained in the captions. Specifically, we generate responses for both the known and unknown questions conditioned on the original and the rewritten descriptions. Information is considered lost if the bi-directional NLI p-entail probabilities are both below 0.5. Our results show that 89% of the unknown information was successfully removed, while 90% of the known information was retained.

### v2

@@ Abstract
[…]
Figure 1: KnowAda identifies knowledge gaps of a VLM and adapts the dense caption accordingly. The KnowAda dense captions are better suited for downstream fine-tuning of the VLM.
@@ 1 Introduction
Fine-tuning pretrained multimodal models for generating dense image captions is common in both research and practical applications, such as assisting visually impaired individuals. Recent work has focused on creating high-quality, descriptive captions through human annotations Onoe et al. (2024); Garg et al. (2024); Deitke et al. (2024) and synthetic generations from models like GPT-4 Chen et al. (2023a, 2024a) and Gemini Singla et al. (2024), including extensions that integrate expert models for enhanced detail Li et al. (2024). These datasets enable the fine-tuning of models to create detailed descriptions in specific styles, distinguishing them from the zero-shot capabilities of pretrained models. However, smaller multimodal models (e.g., up to 7 billion parameters), which are essential for real-time applications, frequently face challenges in capturing fine-grained visual details during fine-tuning, resulting in hallucinations.
@@
Recent works on large language models (LLMs) have shown that fine-tuning primarily adapts pre-existing factual knowledge for specific tasks, with the majority of this knowledge being encoded during the pretraining phase Geva et al. (2020); Meng et al. (2022); Roberts et al. (2020). Rather than acquiring new information, fine-tuning typically activates and refines pretrained knowledge, which remains largely stable throughout the process Zhou et al. (2024b). Furthermore, Gekhman et al. (2024) demonstrated that fine-tuning on content not grounded in a model’s pre-existing factual-knowledge can lead to an increase in hallucinations. Similarly, Yu et al. (2024) showed that attempting to distill GPT4V into a smaller, less capable VLM significantly increases hallucinations. Motivated by these findings, we hypothesize that one possible cause of increased hallucinations is the excessive visual complexity of captions relative to the model’s pretrained capabilities. We propose a method to mitigate this effect.
@@
Evaluating dense image captions requires attention to two critical factors: descriptiveness, which captures the image’s details, and hallucination rate, which measures factual accuracy. Traditional metrics, such as those based on similarity to reference captions Papineni et al. (2002); Banerjee and Lavie (2005); Zhang et al. (2019); Reimers (2019) or CLIP-based alignment Sarto et al. (2023); Radford et al. (2021), often fall short when applied to long captions. These approaches penalize valid variations in phrasing and fail to distinguish between factual accuracy and token overlap or semantic similarity, rendering them inadequate for identifying hallucinations in detailed captions. Existing hallucination metrics are similarly limited, focusing primarily on short captions or object-level errors Rohrbach et al. (2018); Li et al. (2023c); Ben-Kish et al. (2023), or relying on other VLMs Jing et al. (2023); Liu et al. (2023), which themselves can hallucinate during the verification process. To address these limitations, we propose Decomposed NLI (DNLI), a novel evaluation framework that breaks captions into propositions and assesses their entailment with the detailed ground truth description. DNLI offers a more reliable measure of both descriptiveness and accuracy, demonstrating strong alignment with human judgments.
@@
Figure 3: Dense captioning descriptiveness precision-recall results for LLaVA-7B fine-tuned with DOCCI captions, adapted using different methods. “Trimmed” refers to naive removal of sentences, while “Gemini” involves prompting Gemini to simplify the caption by removing difficult details of varying degrees. KnowAda consistently achieves better precision-recall balance. Original captions corresponds to KnowAda with a threshold of $T=100\%$, where no information is classified as unknown.
@@ 2 KnowAda
We begin by introducing our caption adaptation method for dense captioning datasets. KnowAda comprises three stages, as shown in Figure 2. First, we use an LLM to generate visual questions from each dense caption. Next, these questions are employed to probe the VLM’s pretrained visual knowledge and identify parts of the image caption the model struggles with. Finally, an LLM adapts the image descriptions by editing out the unknown parts. Below, we elaborate on each step of the pipeline.
@@ 2.1 VLM Knowledge Probing
4. Calculate the difficulty of each question $q_{i}$ based on the ratio of incorrect to incorrect and correct answers. The difficulty score $df_{i}$ for question $q_{i}$ is given by:
$df_{i}=\frac{|I_{i}|}{|C_{i}|+|I_{i}|}$
@@
For example, consider the two questions shown in Figure 2 (ii). If the model correctly answered ‘‘Monorail’’ for the question ‘‘What kind of vehicle is visible in the image?’’ in 6 out of 10 sampled instances, with 4 incorrect answers, the difficulty for this question is calculated as $df=\frac{4}{10}$ For the second question, ‘‘How many limousines are parked underneath the track?’’, the model correctly answered ‘‘3’’ only 4 out of 10 times, resulting in a difficulty of $df=\frac{6}{10}$. With a threshold of $T=50\%$, the first question is classified as known, while the second as unknown. However, if the threshold is raised to $T=70\%$, both questions would be considered unknown.
For example, consider the two visual questions shown in Figure 2 (ii) with the following prediction accuracies:
  • For the question ‘‘What kind of vehicle is visible in the image?’’, the model correctly answers ‘‘Monorail’’ in $6/10$ instances, resulting in $df=4/10=0.4$.
  • For the question ‘‘How many limousines are parked underneath the track?’’, the model correctly answers ‘‘three’’ in $4/10$ instances, yielding $df=6/10=0.6$.
With a difficulty threshold $T$, the classification of these questions changes:
  1. 1.
    For $T=50\%$:
    • The first question is Known (since $df=0.4<0.5$)
    • The first question is Unknown (since $df=0.6>0.5$)
  2. 2.
    For $T=70\%$:
    • The first question is Unknown (since $df=0.4<0.7$)
    • The second question is Unknown (since $df=0.6<0.7$)
@@
2.2 KnowAda Data Analysis
We start by examining the characteristics of the dense captioning datasets, the impact of KnowAda on the adapted captions, and the types of questions utilized for visual probing within the KnowAda pipeline.
@@ 2.2 KnowAda Data Analysis
Visual questions category distribution. To verify that KnowAda generates diverse types of visual questions, we use LLaMa-3-70B Dubey et al. (2024), quantized to 4 bits, to classify 12,422 questions generated for 1,000 images from the DOCCI test set into categories defined by SeedBench Li et al. (2023a). For the classification task, we provide one-few shot examples from each SeedBench category. The resulting distribution of questions is presented in Figure 5. As observed, the diversity in the question types suggests that KnowAda performs knowledge probing across a wide range of visual tasks.
@@ 3 Decomposed NLI Evaluation
Next, we introduce DNLI, a proposition-decomposition evaluation framework for dense-captioning. It evaluates the quality of a generated dense image caption using two criteria: descriptiveness, which measures accuracy and detail, and contradiction, which measures how much the caption contradicts the ground truth. Inspired by prior work on paragraph summarization evaluation Ernst et al. (2021); Zhang and Bansal (2021); Ernst et al. (2021) and retrieval Chen et al. (2023b), DNLI assesses descriptiveness and consistency through proposition extraction, as outlined below and shown in Figure 4. Additional qualitative examples are in the appendix.
Propositional decomposition. Given a generated image description, we use a Gemini to decompose it into a set of atomic propositions, capturing individual, verifiable claims. This enables fine-grained evaluation. To avoid duplicates, we instruct the model to include only unique propositions.
Natural Language Inference (NLI) analysis. We next evaluate the entailment of propositions generated from the image using both our automatic method, where Gemini is used to compute textual entailment, and through human annotators who directly assess visual entailment:
@@
Note that the neutral labels are discarded in the automatic evaluation, as they could represent either subjective propositions (e.g., "vibrant atmosphere") or visual claims not described in the original captions. These labels could be either entailed or contradicted, but it is unclear which.
Figure 5: Distribution of visual question categories generated from captions in the first stage of KnowAda, classified using SeedBench Li et al. (2023a) definitions.
@@
Table 2: Statistics of the original and KnowAda adapted captions with $T=20\%$ over the human annotated and synthetic version of DOCCI captions. $\bar{C}_{o}$ and $\bar{C}_{r}$ denotes the mean number of words in the original caption and caption after applying KnowAda, and $\bar{Q}_{unk}$ denotes the mean number of questions unknown to the model.
@@
Method Contradiction $\downarrow$ Descriptiveness $\uparrow$ # Words
@@
KnowAda Random 34.7 17.2 57.9 23.9 55
KnowAda 33.7 16.7 58.9 25.8 55
@@ 4 Experimental Settings
  • Caption Trimming: A progressive method that removes varying numbers of sentences to simplify captions.
  • Gemini Simplification: An approach prompting Gemini to remove difficult details from captions to varying degrees.
@@ 5.3 Textual Entailment is Highly Correlated to Human Annotated Visual Entailment
Note that the results from our human annotation process showed an average majority agreement of 82.6% for the “Contradicted” labels and 89.7% for the “Entailed” labels, indicating a high level of consensus among annotators.
5.4 Necessity of Removing Unknown Information
@@ 7.1 Dense Image Captioning
Traditional image captioning datasets like COCO Chen et al. (2015) often feature brief captions with limited visual detail. Recently, interest has grown in creating longer and more complex captions. For example, Onoe et al. 2024 and Garg et al., 2024 developed a dataset of 15K human-annotated, richly descriptive captions. Additionally, Chen et al., 2023a and Singla et al., 2024 employed GPT-4 and Gemini to generate detailed captions. We show that these datasets are more effective for fine-tuning when adapted to the specific model being trained.
7.2 Training On Unknown Information
@@
Several works examine the relationship between unknown information and downstream model performance. Gekhman et al. (2024) show that fine-tuning LLMs on low-confidence examples encourages hallucinations, and suggest that fine-tuning should not introduce new knowledge, but only teach the model to make use of existing knowlegde. Zhang et al. (2024b) suggest optimizing for higher-confidence responses via Direct Preference Optimization (DPO), where the confidence of each response is estimated via self-evaluation. Piché et al. (2024) train a utility function that encourages predicting only high-certainty responses. The function is trained over synthetic data, which is collected iteratively, and is composed of fused high-confidence predicted facts. Additionally, Xu et al. (2024) introduce a refusal mechanism which encourages the model to reject questions that do not align with its existing knowledge. We demonstrate that training a model on questions it is fundamentally lacking the ability to answer, rather than solely on factual knowledge it lacks—particularly in the multimodal domain—results in similar performance degradation. Moreover, our method only rejects the unknown parts of a data sample, while maintaining the information that is useful for training, thus avoiding throwing out useful training samples which only include some unknown parts.
7.3 Data Curation in Fine-Tuning
@@
Recent research has focused on refining the fine-tuning process for pretrained models, especially through instruction tuning. Zhou et al. (2024b) demonstrated that fine-tuning on a small, high-quality instruction dataset can yield superior results. Li et al. (2023b) explored training a language model on a subset of instruction data, using loss discrepancies to inform filtering strategies. However, this approach doesn’t extend to image captioning, where no instructions are available. Lin et al. (2024) proposed reducing noisy token influence by assigning lower weights, while Chen et al. (2024b) trained a network to filter out low-loss instruction samples. In contrast, our approach adapts specific parts of captions misaligned with the model’s capabilities, rather than filtering or reweighing the entire caption. Our strategy leverages visual knowledge and emphasizes sequences and conceptual coherence, moving beyond token-level adjustments.
@@ 8 Limitations
Secondly, while we show that KnowAda achieves better performance in balancing the descriptiveness-hallucination trade-off, it does not fully resolve this issue. Reducing contradictions to the ground truth often decreases descriptiveness recall (but not precision).
@@
Ethics Statement
This work focuses on measuring and mitigating hallucinations in visual-language models (VLMs). As such, it is expected to increase the reliability of VLMs and the ability to measure their performance, which is important when using them in real-world systems. This is expected to have a positive impact on the use of VLMs in society. However, we recognize that the foundation models used in the KnowAda construction and evaluation pipeline could propagate biases. We anticipate further research into such biases before relying on our work beyond the research environment. The human annotation study performed in this work received the required IRB approval from our institution’s ethics committee.
Acknowledgments
This work was partially supported by Google and the TAU Center for Artificial Intelligence and Data Science. The authors thank Morris Alper and Nimrod Shabtay for their valuable feedback and assistance.
@@ Appendix A Appendix
In this appendix, we describe the implementation details in Appendix A.1, qualitative examples in Appendix A.2, additional human annotations details in Appendix A.3 and additional experiments in Appendix A.4.
@@ A.1 Implementation Details
Proposition Entailment Evaluation Implementation Details  We use gemini-1.5-flash-001 both for the proposition extraction, as well as the textual entailment task. We constrain the output to be in JSON format, following the proposition extraction prompt shown in Figure 17 and textual entailment prompt in Figure 18.
@@ A.4 Additional Experiments
DNLI Evaluation with a Smaller Model To justify the use of Gemini in the question difficulty evaluation stage, we compared its performance against Gemma-2B by using both models as judges on 1,000 question-answer pairs from the DOCCI test set. The results showed a 75% agreement between the two models. To further analyze discrepancies, we manually inspected 20 random samples where the judgments differed, finding that in 18 out of 20 cases, Gemini provided the accurate assessment. These findings highlight a trade-off between model size and judgment accuracy. While Gemma-2B is a more lightweight alternative, we prioritized the superior accuracy and consistency offered by Gemini for our evaluations.